# Manydepth2: Motion-Aware Self-Supervised Monocular Depth Estimation in Dynamic Scenes

### v2

@@ III-C New Static Reference Frame (Rewrite and Check)
where $\epsilon$ is the threshold for distinguishing moving parts. The motion mask in the target mask can be used to generate the motion mask $\mathcal{M}m^{r}$ in the reference mask, utilizing $f{r\leftarrow t}$. With $\mathcal{M}_{m}^{r}$, it becomes possible to reconstruct a new static reference image $\hat{I_{r}}$ by eliminating the impact of dynamic objects. This process can be formulated as follows:
@@ IV Experiments
TABLE I: Self-supervised depth estimation results on the KITTI 2015. We evaluate the performance of algorithms utilizing either stereo or monocular information during the training phase. Metrics denoted by Color have a lower optimal value, while metrics denoted by Color have a higher optimal value. The optimal results in each subsection are denoted in bold. Our method outperforms all prior approaches in most metrics across all subsections, irrespective of whether the baselines utilize multiple frames during testing. M: Multi-frames. S: Motion Segmentation. F: Optical Flow.
@@ IV-A Monocular Depth Estimation
KITTI-2015: The Tab. I presents a comparison between the performance of Manydepth2 and state-of-the-art algorithms. Overall, Manydepth2 exhibits superior performance compared to other methods, irrespective of whether monocular or multiview images are used. In particular, Manydepth2 demonstrates a significant performance advantage over Manydepth, which also employs multiview images and a cost volume structure to estimate depth. Manydepth2 outperforms Manydepth by a substantial margin of $7.2\%$ in terms of Abs Rel. In comparison to Dynamicdepth, which employs semantic motion segmentation maps to leverage high-level computer vision information for precise depth estimation, Manydepth2 achieves a superior performance of $5.3\%$ in terms of the Abs Rel, without the inclusion of any domain-specific information. Quantitative results are shown in Fig. 4.
Cityscapes: In contrast to the KITTI dataset, the Cityscapes dataset features a higher percentage of dynamic objects. During the training phase on the Cityscapes dataset, the input resolution of Manydepth2 is set to $128\times 416$. The lower section of Tab. II indicates a substantial performance advantage of Manydepth2 over its competing methods. Remarkably, Manydepth2 exhibits superior performance over ManyDepth, with a margin of approximately $15.0\%$ in the absolute relative error metric. Through the visualization of error maps generated by computing the differences between predicted depth maps and their corresponding ground truth depth maps, Fig. 1 facilitates a qualitative analysis of depth predictions. These observations indicate that Manydepth2 offers enhanced accuracy for depth estimation of both dynamic foreground and static background. Additionally, the incorporation of a pre-trained optical flow estimation model with low computational requirements has resulted in a minimal increase in parameter size and running time for Manydepth2. Quantitative results are shown in Fig. 1. Tab. III showcases the depth errors within dynamic object regions by utilizing the ground truth motion segmentation maps provided in [16], which corresponds to the qualitative results presented in Fig. 1.
@@
w/o Motion-Guided Volume 0.094 0.676 4.289
w/o Attention for Final Volume 0.093 0.648 4.338
w Normal HRNet 0.096 0.746 4.349
Manydepth2 0.091 0.650 4.263
@@ IV-C Abaltion Study
In our research, we meticulously carry out ablation studies using the Cityscapes dataset. These studies are designed to dissect and showcase how effective our proposed modules are, allowing us to validate the specific design choices we’ve made. The detailed outcomes of these experiments are showcased in Table V. Within this table, we provide insights from experiments conducted under diverse conditions and configurations. Specifically, we assess Manydepth2’s performance with variations such as the Prior Depth Net, Manydepth2 configurations without Motion-Guided Volume, without Attention for Final Volume, and with the inclusion of a Normal HRNet. A key observation from Table V is the pronounced positive impact that integrating the Attention-Based HRNet has on enhancing Manydepth2’s performance. Furthermore, our findings underscore that the incorporation of both the Motion-Guided Volume and Attention for the Final Volume amplifies the capabilities of Manydepth2 even further, leading to improved outcomes.

### v3

@@ III-C New Static Reference Frame (Rewrite and Check)
where $\epsilon$ is the threshold for distinguishing moving parts. The motion mask in the target mask can be used to generate the motion mask $\mathcal{M}_{m}^{r}$ in the reference mask, utilizing $f_{r\leftarrow t}$. With $\mathcal{M}_{m}^{r}$, it becomes possible to reconstruct a new static reference image $\hat{I_{r}}$ by eliminating the impact of dynamic objects. This process can be formulated as follows:
@@ IV Experiments
TABLE I: Self-supervised depth estimation results on the KITTI 2015. We evaluate the performance of algorithms utilizing either stereo or monocular information during the training phase. Metrics denoted by Color have a lower optimal value, while metrics denoted by Color have a higher optimal value. The optimal results in each subsection are denoted in bold. Our method outperforms all prior approaches in most metrics across all subsections, irrespective of whether the baselines utilize multiple frames during testing. M: Multi-frames. S: Motion Segmentation. F: Optical Flow. Manydepth2-NF refers to Manydepth2 without optical flow.
@@ IV-A Monocular Depth Estimation
KITTI-2015: The Tab. I presents a comparison between the performance of Manydepth2 and state-of-the-art algorithms. Overall, Manydepth2 exhibits superior performance compared to other methods, irrespective of whether monocular or multiview images are used. In particular, Manydepth2 demonstrates a significant performance advantage over Manydepth, which also employs multiview images and a cost volume structure to estimate depth. Manydepth2 outperforms Manydepth by a substantial margin of $7.2\%$ in terms of Abs Rel. In comparison to Dynamicdepth, which employs semantic motion segmentation maps to leverage high-level computer vision information for precise depth estimation, Manydepth2 achieves a superior performance of $5.3\%$ in terms of the Abs Rel, without the inclusion of any domain-specific information. Quantitative results are shown in Fig. 4. We also present the results for Manydepth2-NF, a version of Manydepth2 that does not use optical flow or the motion-guided cost volume. We found that our attention-based network outperformed Manydepth by $4.1\%$.
Cityscapes: In contrast to the KITTI dataset, the Cityscapes dataset features a higher percentage of dynamic objects. During the training phase on the Cityscapes dataset, the input resolution of Manydepth2 is set to $128\times 416$. The lower section of Tab. II indicates a substantial performance advantage of Manydepth2 over its competing methods. Remarkably, Manydepth2 exhibits superior performance over ManyDepth, with a margin of approximately $15.0\%$ in the absolute relative error metric. Through the visualization of error maps generated by computing the differences between predicted depth maps and their corresponding ground truth depth maps, Fig. 1 facilitates a qualitative analysis of depth predictions. These observations indicate that Manydepth2 offers enhanced accuracy for depth estimation of both dynamic foreground and static background. Additionally, the incorporation of a pre-trained optical flow estimation model with low computational requirements has resulted in a minimal increase in parameter size and running time for Manydepth2. Quantitative results are shown in Fig. 1. Tab. III showcases the depth errors within dynamic object regions by utilizing the ground truth motion segmentation maps provided in [16], which corresponds to the qualitative results in Fig. 1.
@@
w/o Attention for Volume 0.100 0.709 4.298
Manydepth2-NF 0.094 0.676 4.246
Manydepth2 0.091 0.649 4.232
@@ IV-C Abaltion Study
In our research, we meticulously carry out ablation studies using the Cityscapes dataset. These studies are designed to dissect and showcase how effective our proposed modules are, allowing us to validate the specific design choices we’ve made. The detailed outcomes of these experiments are showcased in Table V. Within this table, we provide insights from experiments conducted under diverse conditions and configurations. Specifically, we assess Manydepth2’s performance with variations such as the Prior Depth Net, Manydepth2 configurations without Motion-Guided Volume and without Attention for Final Volume. A key observation from Table V is the pronounced positive impact that integrating the Attention-Based HRNet has on enhancing Manydepth2’s performance. Furthermore, our findings underscore that the incorporation of both the Motion-Guided Volume and Attention for the Final Volume amplifies the capabilities of Manydepth2 even further, leading to improved outcomes.